# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v3

@@ 5.1 Gradient domination with $\zeta=1$
Given an MAB problem with $K$ arms and known deterministic rewards $r\in[0,1]^{K}$, consider the class of softmax policies $\pi_{\theta}\in\Delta_{K}$ parameterized by $\theta\in\mathbb{R}^{K}$ s.t. $\pi_{\theta}(a)=\frac{\exp(\theta(a))}{\sum_{a^{\prime}}\exp(\theta(a^{\prime}))}$. The loss corresponding to the bandit problem is given by: $f(\theta)=r(a^{*})-\langle\pi_{\theta},r\rangle$, where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=72$Assn. 3 with $\nu=24$ and $\omega=0$ and Assn. 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$.
@@
The above result is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms designed for this specific problem, including GD with a specific line-search that requires the knowledge of $f^{*}$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Kakade & Langford, 2002; Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024). From a practical perspective, Lu et al. (2024, Figure 1) empirically demonstrate the linear convergence of GD-LS on tabular Markov decision processes, and hence, Cor. 6 substantiates their results theoretically.
Two-layer neural networks: We note that Assn. 1, 2 and 3 and Assn. 4 with $\zeta=1$ are also satisfied when using the (i) exponential loss to train (ii) two-layer neural networks with a smoothed leaky-ReLU non-linearity and (iii) assuming that the training data is linearly separable (Taheri & Thrampoulidis, 2023, Lemmas 3,5). In this setting, Theorem 1 in Taheri & Thrampoulidis (2023) shows that normalized GD can result in linear convergence for the resulting non-convex objective. Since this problem also satisfies the required assumptions for Cor. 3, GD-LS also converges linearly and unlike normalized GD, it does not require knowing problem-dependent constants.
@@ 6 SGD with Stochastic Line-search
In this section, we analyze the convergence of stochastic gradient descent (SGD) (Robbins & Monro, 1951) with a stochastic variant of the Armijo line-search (referred to as SGD-SLS) proposed in Vaswani et al. (2019b).
@@
In order to analyze the convergence of SGD-SLS, we define $f_{i}^{*}:=\min f_{i}(\theta)$ as the minimum of function $i$ in the finite-sum and $\chi^{2}(\theta^{*}):=\mathbb{E}_{i}[f_{i}(\theta^{*})-f_{i}^{*}]$ as the measure of the stochasticity at the optimum (Loizou et al., 2021). In particular, if $\chi^{2}=0$, then each $f_{i}$ is minimized at $\theta^{*}$ implying that $\nabla f_{i}(\theta^{*})=0$. This special case is referred to as the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) and is useful in practical machine learning; for example, it is approximately satisfied by over-parameterized neural networks (Zhang et al., 2017) or non-parametric regression (Liang & Rakhlin, 2018; Belkin et al., 2019). Furthermore, logistic regression on linearly separable data is an example of a smooth convex loss that satisfies the interpolation condition and is the main motivation for the subsequent analysis.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,1\right)$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ requires $T=O\left(\frac{1}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right)$ iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq\frac{5\epsilon}{2}$.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{C^{\prime}}{8}\right)$ where $C^{\prime}=O(1)$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ requires $T=O\left(\frac{n}{\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}\right)$ iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq\frac{5\epsilon}{2}$.
@@ 7 Conclusion
We believe that analyzing GD-LS for a broader class of non-convex functions, and characterizing the advantage of using Armijo-LS for other algorithms (such as Nesterov accelerated gradient) are important future directions. We also plan to investigate whether other adaptive step-size schemes such as AdaGrad (Duchi et al., 2011) or Adam (Kingma & Ba, 2014) can also provably adapt to the non-uniform smoothness and result in faster convergence rates.
@@ Acknowledgments
We thank Damien Scieur, Mark Schmidt, Curtis Fox, Michael Lu and Siyi Meng for helpful discussions and feedback, and Anh Dang for help with the initial experiments. This research was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2022-04816.
@@
Case 1: If $L_{1}=0$Assn. 2 is equivalent to the standard $L_{0}$-uniform smoothness condition. In this case, we can follow the standard analysis of GD-LS (Nocedal & Wright, 2006) and conclude that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L_{0}}\right\}\geq\min\left\{\eta_{\max},\frac{(1-c)}{3\,L_{0}}\right\}$.
@@
We now follow a proof similar to that of Axiotis & Sviridenko (2023, Theorem 5.2) and derive a linear rate of convergence. From the theorem assumption, we know that $\left\|\nabla f({\theta_{t}})\right\|_{2}^{2}\geq\frac{[f({\theta_{t}})-f^{*}]^{2}}{R}$. Combining these relations,
@@ C.2 Helper Lemmas
For the $\epsilon$ and $T$ defined in the theorem statement, define the event $G=\{\inf_{t\in[T]}[f(\theta_{t})]\leq\epsilon\}$. Using the law of total expectation,
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $=\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G\right]\,\Pr[G]+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]\,\Pr[G^{c}]$ (28)
$\leq\epsilon+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]$ (By definition of $G$)
Hence, we need to bound $\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ conditioned on the event $G^{c}=\{\inf_{t\in[T]}f(\theta_{t})\geq\epsilon\}$.
Conditioned on $G^{c}$: For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$. Using the standard guarantee for Armijo line-search, we know that $\eta_{\max}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L}\right\}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5,
@@
Taking an expectation over the randomness in iteration $t$, conditioned on the past iterates,
$\implies\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{f(\theta_{t})}{3}+\frac{2f(u)}{\epsilon}$ (Since $u$ is independent of the randomness at iteration $t$)
@@
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{f(\theta_{t})}{3}+\frac{\epsilon}{6}$
From the conditioning on $G^{c}$, we know that $f(\theta_{t})\geq\epsilon$. Hence,
$\implies\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\epsilon}{3}+\frac{\epsilon}{6}=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\epsilon}{6}$
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]-\frac{\epsilon}{6}$
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\epsilon\,T}{6}$ (Summing from $t=0$ to T1T-1)
$\leq\beta^{2}-\underbrace{\frac{\epsilon}{6}}_{:=C}\,T$ (Since $\theta_{0}=0$ and $\left\|u\right\|=\beta$)
$T\geq\frac{\beta^{2}}{C}$ $=\frac{6}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{12}{\epsilon^{2}}\right)\right]^{2}$
Since the above calculations were conditioned on $G^{c}$, we can conclude that after $T=\frac{6}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{12}{\epsilon^{2}}\right)\right]^{2}$ iterations, $\mathbb{E}[\inf_{t\in[T]}f(\theta_{T})|G^{c}]\leq\frac{3\epsilon}{2}$. Hence, from Eq. 29
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $\leq\frac{5\epsilon}{2}$
@@
For the $\epsilon$ and $T$ defined in the theorem statement, define the event $G=\{\inf_{t\in[T]}[f(\theta_{t})]\leq\epsilon\}$. Using the law of total expectation,
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $=\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G\right]\,\Pr[G]+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]\,\Pr[G^{c}]$ (29)
$\leq\epsilon+\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})|G^{c}\right]$ (By definition of $G$)
Hence, we need to bound $\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ conditioned on the event $G^{c}=\{\inf_{t\in[T]}f(\theta_{t})\geq\epsilon\}$.
Conditioned on $G^{c}$: Note that from Lemma 1, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{C^{\prime}}{{f_{t}}(\theta_{t})}\right\}$ where $C^{\prime}:=\frac{1}{\lambda_{1}}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5, we get that,
@@
$(*)$ $=\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}+\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})<\epsilon\}$
$\implies(*)$ $\geq\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}\geq\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}$
@@
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\geq\epsilon\}+2\eta_{\max}[{f_{t}}(u)]$
Taking an expectation over the randomness in iteration $t$, conditioned on the past iterates,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,\Pr[f_{t}(\theta_{t})\geq\epsilon]+2\eta_{\max}[f(u)]$ (Since $\eta_{\max}$ and $u$ are both independent of the randomness at iteration $t$)
From the conditioning on $G^{c}$, we know that $f(\theta_{t})\geq\epsilon$. Since $f$ is the average of the $f_{i}$ losses, we know that there exists at least one $j$ s.t. $f_{j}(\theta_{t})\geq\epsilon$. Hence, conditioned on the past iterates, $\Pr[f_{t}(\theta_{t})]\geq\epsilon\geq\frac{1}{n}$. Combining this lower bound with the above inequality, we get that,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2n}+2\eta_{\max}[f(u)]$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{2n}+\frac{2\,f(u)}{\epsilon}$ (Setting $\eta_{\max}=\frac{1}{\epsilon}$)

### v4

@@ 5.1 Gradient domination with $\zeta=1$
Given an MAB problem with $K$ arms and known deterministic rewards $r\in[0,1]^{K}$, consider the class of softmax policies $\pi_{\theta}\in\Delta_{K}$ parameterized by $\theta\in\mathbb{R}^{K}$ s.t. $\pi_{\theta}(a)=\frac{\exp(\theta(a))}{\sum_{a^{\prime}}\exp(\theta(a^{\prime}))}$. The loss corresponding to the bandit problem is given by: $f(\theta)=r(a^{*})-\langle\pi_{\theta},r\rangle$, where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=72$, Assn. 3 with $\nu=24$ and $\omega=0$ and Assn. 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$.
@@
The above result is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms designed for this specific problem, including GD with a specific line-search that requires the knowledge of $f^{*}$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Kakade and Langford, 2002; Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024). From a practical perspective, Lu et al. (2024, Figure 1) empirically demonstrate the linear convergence of GD-LS on tabular Markov decision processes, and hence, Cor. 6 substantiates their results theoretically.
Two-layer neural networks: We note that Assn. 1, 2 and 3 and Assn. 4 with $\zeta=1$ are also satisfied when using the (i) exponential loss to train (ii) two-layer neural networks with a smoothed leaky-ReLU non-linearity and (iii) assuming that the training data is linearly separable (Taheri and Thrampoulidis, 2023, Lemmas 3,5). In this setting, Theorem 1 in Taheri and Thrampoulidis (2023) shows that normalized GD can result in linear convergence for the resulting non-convex objective. Since this problem also satisfies the required assumptions for Cor. 3, GD-LS also converges linearly and unlike normalized GD, it does not require knowing problem-dependent constants.
@@ 6 SGD with Stochastic Line-search
In this section, we analyze the convergence of stochastic gradient descent (SGD) (Robbins and Monro, 1951) with a stochastic variant of the Armijo line-search (referred to as SGD-SLS) proposed in Vaswani et al. (2019b).
@@
In order to analyze the convergence of SGD-SLS, we define $f_{i}^{*}:=\min f_{i}(\theta)$ as the minimum of function $i$ in the finite-sum and $\chi^{2}(\theta^{*}):=\mathbb{E}_{i}[f_{i}(\theta^{*})-f_{i}^{*}]$ as the measure of the stochasticity at the optimum (Loizou et al., 2021). In particular, if $\chi^{2}=0$, then each $f_{i}$ is minimized at $\theta^{*}$ implying that $\nabla f_{i}(\theta^{*})=0$. This special case is referred to as the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt and Roux, 2013) and is useful in practical machine learning; for example, it is approximately satisfied by over-parameterized neural networks (Zhang et al., 2017) or non-parametric regression (Liang and Rakhlin, 2018; Belkin et al., 2019). Furthermore, logistic regression on linearly separable data is an example of a smooth convex loss that satisfies the interpolation condition and is the main motivation for the subsequent analysis.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,1\right)$, if $\tau_{\epsilon}:=\inf\{t:f(\theta_{t})\leq\epsilon\}$ is the hitting time, then SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ ensures that, $\mathbb{E}[\tau_{\epsilon}]=O\left(\frac{1}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right).$
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,C^{\prime}\right)$ with $C^{\prime}=O(1)$, if $\tau_{\epsilon}:=\inf\{t:f(\theta_{t})\leq\epsilon\}$ is the hitting time, then SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ ensures that, $\mathbb{E}[\tau]=O\left(\frac{n}{\gamma^{2}}\,\left[\ln(\frac{n}{\epsilon^{2}})\right]^{2}\right).$
@@ 7 Conclusion
We believe that analyzing GD-LS for a broader class of non-convex functions, and characterizing the advantage of using Armijo-LS for other algorithms (such as Nesterov accelerated gradient) are important future directions. We also plan to investigate whether other adaptive step-size schemes such as AdaGrad (Duchi et al., 2011) or Adam (Kingma and Ba, 2014) can also provably adapt to the non-uniform smoothness and result in faster convergence rates.
@@ Acknowledgments
We thank Anant Raj for pointing out and helping to resolve a mistake in a previous version of the paper. We thank Damien Scieur, Mark Schmidt, Curtis Fox, Michael Lu and Siyi Meng for helpful discussions and feedback, and Anh Dang for help with the initial experiments. This research was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2022-04816.
@@
Case 1: If $L_{1}=0$, Assn. 2 is equivalent to the standard $L_{0}$-uniform smoothness condition. In this case, we can follow the standard analysis of GD-LS (Nocedal and Wright, 2006) and conclude that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L_{0}}\right\}\geq\min\left\{\eta_{\max},\frac{(1-c)}{3\,L_{0}}\right\}$.
@@
We now follow a proof similar to that of Axiotis and Sviridenko (2023, Theorem 5.2) and derive a linear rate of convergence. From the theorem assumption, we know that $\left\|\nabla f({\theta_{t}})\right\|_{2}^{2}\geq\frac{[f({\theta_{t}})-f^{*}]^{2}}{R}$. Combining these relations,
@@ C.2 Helper Lemmas
If $i_{t}$ indicates the index of the function selected randomly at iteration $t$, define ${\mathcal{F}}_{t}=\sigma\{i_{0},i_{1},\ldots,i_{t-1}\}$ as the filtration for $t\geq 0$. We define $\tau:=\inf\{t:f(\theta_{t})\leq\epsilon\}$ as the (random) hitting time. Note that ${\mathcal{F}}_{t}$ contains all the randomness until iteration $t$, and that the event $\{t<\tau\}=\{f(\theta_{0})>\epsilon\}\cap\{f(\theta_{1})>\epsilon\}\cap\ldots\cap\{f(\theta_{t-1})>\epsilon\}$ only depends on the randomness in iterations 0 to t1t-1.
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$. Using the standard guarantee for Armijo line-search, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{2\,(1-c)}{L}\right\}$ and ${\eta_{t}}\leq\eta_{\max}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5,
@@
Taking expectation w.r.t. randomness at iteration $t$ conditioned on ${\mathcal{F}}_{t}$, we have
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{3}\,f(\theta_{t})+\frac{2f(u)}{\epsilon}\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\epsilon}{3}\,{\mathcal{I}}\{f(\theta_{t})>\epsilon\}+\frac{2f(u)}{\epsilon}$
By multiplying both-sides by ${\mathcal{I}}\{t<\tau\}$ we get,
$\implies\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]\,{\mathcal{I}}\{t<\tau\}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}-\frac{\epsilon}{3}\,{\mathcal{I}}\{f(\theta_{t})>\epsilon\cap t<\tau\}+\frac{2f(u)}{\epsilon}\,{\mathcal{I}}\{t<\tau\}$
Now we take expectation w.r.t. all randomness from iteration 0 to t1t-1,
$\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]$ $\leq\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\epsilon}{3}\,\mathbb{E}_{0:t-1}[{\mathcal{I}}\{f(\theta_{t})>\epsilon\cap t<\tau\}]+\frac{2f(u)}{\epsilon}\,\mathbb{E}_{0:t-1}[{\mathcal{I}}\{t<\tau\}]$
$=\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\epsilon}{3}\,\Pr[\{f(\theta_{t})>\epsilon\}\cap\{t<\tau\}]+\frac{2f(u)}{\epsilon}\,\Pr[\{t<\tau\}]$
$\implies\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]$ $\leq\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\epsilon}{3}\,\Pr[t<\tau]+\frac{2f(u)}{\epsilon}\,\Pr[\{t<\tau\}]$ (Using that $\Pr[\{f(\theta_{t})>\epsilon\}|\{t<\tau\}]=1$)
@@
$\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]$ $\leq E_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\epsilon}{3}\,\Pr\{t<\tau\}\,+\frac{\epsilon}{6}\,\Pr\{t<\tau\}$
$\implies\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t+1<\tau\}]$ $\leq E_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\epsilon}{6}\,\Pr\{t<\tau\}$ (Using that ${\mathcal{I}}\{t+1<\tau\}<{\mathcal{I}}\{t<\tau\}$)
By rearranging, summing both-sides, telescoping we have,
$\frac{\epsilon}{6}\,\sum_{t=0}^{\infty}\Pr\{t<\tau\}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}\,{\mathcal{I}}\{0<\tau\}\leq\left\|\theta_{0}-u\right\|_{2}^{2}$
$\implies\sum_{t=0}^{\infty}\Pr\{t<\tau\}$ $\leq O\left(\frac{1}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right)$ (By setting $\theta_{0}=0$ and since $\beta=\frac{1}{\gamma}\ln(\frac{12}{\epsilon^{2}})$)
Since $\tau=\sum_{t=0}^{\infty}{\mathcal{I}}\{t<\tau\}$,
$\mathbb{E}[\tau]=\mathbb{E}\left[\sum_{t=0}^{\infty}{\mathcal{I}}\{t<\tau\}\right]\leq\sum_{t=0}^{\infty}\Pr\{t<\tau\}$ (Using Fubini’s theorem)
By combining the above results, we get that $\mathbb{E}[\tau]=O(\frac{1}{\epsilon\,\gamma^{2}}\,\left[\ln(\frac{1}{\epsilon^{2}})\right]^{2})$. ∎
@@
The proof is similar to that of Thm. 4. In particular, if $i_{t}$ indicates the index of the function selected randomly at iteration $t$, define ${\mathcal{F}}_{t}=\sigma\{i_{0},i_{1},\ldots,i_{t-1}\}$ as the filtration for $t\geq 0$. We define $\tau:=\inf\{t:f(\theta_{t})\leq\epsilon\}$ as the (random) hitting time. Note that ${\mathcal{F}}_{t}$ contains all the randomness until iteration $t$, and that the event $\{t<\tau\}=\{f(\theta_{0})>\epsilon\}\cap\{f(\theta_{1})>\epsilon\}\cap\ldots\cap\{f(\theta_{t-1})>\epsilon\}$ only depends on the randomness in iterations 0 to t1t-1.
Note that from Lemma 1, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{C^{\prime}}{{f_{t}}(\theta_{t})}\right\}$ where $C^{\prime}:=\frac{1}{\lambda_{1}}$ and ${\eta_{t}}\leq\eta_{\max}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5, we get that,
@@
$(*)$ $=\frac{\min\left\{\eta_{\max}\,{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\}+\frac{\min\left\{\eta_{\max}\,{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})\leq\epsilon\}$
$\implies(*)$ $\geq\frac{\min\left\{\eta_{\max}\,{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\}\geq\frac{\min\left\{\eta_{\max}\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\}$
@@
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\,\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\}+2\eta_{\max}[{f_{t}}(u)]$
Taking expectation w.r.t. randomness at iteration $t$ conditioned on ${\mathcal{F}}_{t}$, we have
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}\,\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\}+2\eta_{\max}[f(u)]$ (Since ${\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\}$ does not depend on the randomness at iteration $t$)
By multiplying both-sides by ${\mathcal{I}}\{t<\tau\}$ we get,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]\,{\mathcal{I}}\{t<\tau\}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}-\frac{\min\left\{\eta_{\max}\,\epsilon,C^{\prime}\right\}}{2}\,{\mathcal{I}}\{f_{t}(\theta_{t})>\epsilon\cap t<\tau\}+2\eta_{\max}[f(u)]\,{\mathcal{I}}\{t<\tau\}$
Now we take expectation w.r.t. all randomness from iteration 0 to t1t-1,
$\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]$
$\leq\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\min\left\{\eta_{\max}\,\epsilon,C^{\prime}\right\}}{2}\,\Pr[\{f_{t}(\theta_{t})>\epsilon\cap t<\tau\}]+2\eta_{\max}[f(u)]\Pr[\{t<\tau\}]$
$=\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\min\left\{\eta_{\max}\,\epsilon,C^{\prime}\right\}}{2}\,\Pr[f_{t}(\theta_{t})>\epsilon|\{t<\tau\}]\,\Pr[\{t<\tau\}]+2\eta_{\max}[f(u)]\Pr[\{t<\tau\}]$
From the conditioning on $\{t<\tau\}$, we know that $f(\theta_{t})>\epsilon$. Since $f$ is the average of the $f_{i}$ losses, we know that there exists at least one $j$ s.t. $f_{j}(\theta_{t})>\epsilon$. Hence, $\Pr[f_{t}(\theta_{t})>\epsilon|\{t<\tau\}]\geq\frac{1}{n}$. Combining this lower bound with the above inequality, we get that,
$\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]\leq\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\min\left\{\eta_{\max}\,\epsilon,C^{\prime}\right\}}{2n}\,\Pr[\{t<\tau\}]+2\eta_{\max}[f(u)]\Pr[\{t<\tau\}]$
Setting $\eta_{\max}=\frac{1}{\epsilon}$ and using that ${\mathcal{I}}\{t+1<\tau\}<{\mathcal{I}}\{t<\tau\}$,
$\implies$ $\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t+1<\tau\}]\leq\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\min\left\{1,C^{\prime}\right\}}{2n}\,\Pr[\{t<\tau\}]+\frac{2\,f(u)}{\epsilon}\,\Pr[\{t<\tau\}]$